# LipidBERT: A Lipid Language Model Pre-trained on METiS de novo Lipid Library

### v1

@@ 1 Introduction
In this paper, we present LipidBERT, a model pre-trained on an in silico generated and meticulously screened lipid library constructed via fragment-based generative methods and reinforcement learning (RL). Molecular fragments with both lipid and small molecule features were generated by RL and connected to form ionizable lipids. These were then subjected to a series of knowledge-, molecular dynamics-, and AI-based filters. The resulting library consists of 10 million ionizable lipids, which were subsequently used for pre-training several BERT-like [21] models on different data scales. The models were then fine-tuned for downstream LNP property prediction tasks, using wet-lab labeled data from our experimental team. The fine-tuned prediction model exhibits a remarkably high $R^{2}$ score (over 0.9), demonstrating the power of pre-trained lipid language models on virtual datasets. Apart from language models, we also have graph-based and several distribution-based models, which might be reported in the near future.
@@ 2.3.3 Connecting Atom Prediction - Sequence Classification
Figure 5: Visualization of the predicted connecting atom between head and tail via sequence classification, using the connecting points prediction head in the pre-trained model. The predicted atoms are explicitly marked. The visualization was created using RDKit [22].
@@ 2.4 Fine-Tuning
For each pre-trained model, we use the 768-dimensional embeddings of the [CLS] token to represent the entire sequence. The sequence output [CLS] 768-dimensional embeddings pass through a linear layer and a tanh activation function, followed by an MLP layer to generate the final output. The MLP layer consists of four layers that gradually transform the embeddings from 768 to 128 dimensions, applying the GeLU [26] activation function. A final linear layer is employed to generate the output for regression tasks. In the context below, LipidBERT refers to the model pre-trained with or without secondary tasks that achieves the best performance (highest validation $R^{2}$ value).
@@ 2.4.1 Fine-Tuning Using Our Wet-Lab Experimental Dataset
The data used for fine-tuning originated from our experimental team and includes various properties of lipid nanoparticles (LNPs). We fine-tuned each pre-trained model, which had been trained on different scales of training data, based on various experimental LNP properties. Here we present the fine-tuned results for some of the aforementioned LNP properties as examples. It is important to note that the LNP data used in this study were all generated using consistent ratios of the various components in the formulation. The primary aim of these experiments was to screen and identify the optimal ionizable lipids for use in the subsequent development of cell-, tissue-, or organ-targeting LNPs.
Figure 9: Validation set $R^{2}$ values change along epochs during LipidBERT fine-tuning, concerning various LNP properties.
Various LNP Properties Prediction
First, we tested our LipidBERT model on downstream tasks of sequence regression using wet-lab data of selected various LNP properties, and the results are presented in Figure 9. The figure demonstates that our LipidBERT model exhibits exceptional performance, yielding validation $R^{2}$ values greater than 0.9 in most of the regression tasks, with the exception of toxicity (0.79). The in-vivo overall fluorescence intensity shows the fastest convergence, stabilizing around epoch 75. Additionally, unlike other LNP properties, it demonstrates the most steady increase, whereas others fluctuate significantly more. Moreover, the polymer dispersity index (PDI) also shows a relatively steady increase after 25 epochs. Most of the $R^{2}$ curves fluctuate significantly before 60 epochs, turning from negative to positive values after 60 epochs and then gradually rising. It is important to note that the prediction model performs exceptionally well for ionizable lipids with similar scaffolds to those in our wet-lab data but may provide less accurate property predictions for lipids with entirely new scaffolds. In other words, our LipidBERT model may produce deviated prediction values if the given lipids have structures quite different from our experimental ones. However, the prediction model continues to evolve as the amount and diversity of our experimental lipids grow.
Figure 10: Validation set $R^{2}$ values change along epochs concerning (a) ex-vivo fluorescence intensity in lung, and (b) lipid nanoparticle size. Different curves represent LipidBERT models pre-trained on different scales of training dataset.
@@
The effect of training set sizes is also explored by fine-tuning LipidBERT models pre-trained on various training set sizes using the same experimental data mentioned above. The variation of the validation $R^{2}$ values over training epochs is presented in Figure 10. Maximum validation $R^{2}$ values are selected to evaluate the performance of each model. The models were trained for different numbers of epochs on the two experimental datasets to achieve more stabilized validation $R^{2}$ values. From both figures, it is evident that model performance improves as the size of the training dataset increases. In Figure 10(a), there is a dramatic enhancement in $R^{2}$ from 0.26 with the 0.05M training set to 0.53 with the 0.25M training set, and further to 0.80 with the 1M training set. Increasing the training dataset size from 1M to 3M boosts the $R^{2}$ to 0.89, while expanding the dataset from 3M to 10M yields a smaller increase, with R2 reaching 0.94. In contrast, Figure 10(b) shows that the 0.05M and 0.25M models exhibit negligible differences in predicting lipid nanoparticle size, with $R^{2}$ values of 0.35 and 0.34, respectively. However, there is a significant increase in $R^{2}$ between the 0.25M and 1M models, where the $R^{2}$ jumps from 0.34 to 0.72. Further increases in dataset size eventually yield a 10M model with an $R^{2}$ value of 0.91. These examples demonstrate that increasing the training dataset size is critical to obtaining a fine-tuned model with high $R^{2}$ values, highlighting the importance of large datasets for model accuracy.
Figure 11: Validation set $R^{2}$ values of LipidBERT and various XGBoost models trained on combinations of different descriptors.
@@
We also compared the fine-tuned 10M-LipidBERT model with our previously developed XGBoost [27] model, which is trained using combinations of regular molecular descriptors, Molecular Dynamics (MD) descriptors, and Quantum Mechanics (QM) descriptors. The regular descriptors are generated via RDKit [22], while the MD and QM descriptors are generated from MD and QM simulations, respectively. The MD descriptors are obtained from our fully-automated high-throughput (HTS) lipid bilayer simulations, designed to mimic the lipid movement in realistic LNPs. Given a lipid SMILES, the HTS platform can automatically construct the lipid/bilayer 3D structures, perform MD simulations, and extract the MD descriptors from the MD trajectories without human intervention. Similarly, the QM descriptors are also obtained automatically from our HTS platform. QM calculations are performed on a single ionizable lipid molecule to extract the QM descriptors. All descriptors are combined and used for XGBoost training on the overall fluorescence intensity, and the results are presented in Figure 11. The figure shows that the LipidBERT model exhibits overwhelming advantages over all XGBoost models, achieving an extremely high $R^{2}$ value of 0.98. Among the XGBoost models, the QM descriptors-based model gives the highest $R^{2}$ value of 0.63, indicating that QM simulations can derive better descriptors than regular and MD descriptors.
@@
Apart from LipidBERT, we also trained PhatGPT, our GPT-like [28] lipid generation model using the same dataset as LipidBERT. Despite the fact that PhatGPT is originally designed and trained for generative tasks, we compared the performance of the two language models on downstream sequence regression tasks. It is important to note that the vanilla PhatGPT has different model parameters (such as the number of attention layers and embedding sizes). For the purpose of comparison, we pre-trained another PhatGPT model using the same parameters as LipidBERT. Several LNP properties were used for fine-tuning, but PhatGPT did not perform as expected, yielding validation $R^{2}$ values around 0.01. This outcome further demonstrates the superiority of BERT-like models in sequence regression/classification tasks, while GPT-like models are primarily designed for token generation.
@@ 2.4.2 Fine-Tuning Using the AGILE Public Dataset
We followed the official instructions on AGILE’s github page (https://github.com/bowang-lab/AGILE) to fine-tune the pre-trained AGILE model utilizing its provided dataset, and compared it with our fine-tuned LipidBERT model using the same dataset. Both models were fine-tuned for 100 epochs using the same random seed for an 80%-10%-10% train-test-validation dataset split. The fine-tuned results, including validation loss, Pearson correlation coefficient and $R^{2}$, are given in Figure 12. Generally speaking, our transformer-based LipidBERT model outperforms the graph-based AGILE model in all evaluation metrics. For validation loss (Figure 12(a)), LipidBERT experiences a gradual decline from 30 to 0.4 for the Hela dataset and from 8.5 to 0.2 for the Raw 264.7 dataset, which are lower than the AGILE validation losses (12.8 and 2.1). The Pearson correlation coefficient ranges from -1 to 1, where -1 represents a perfect negative linear relationship and 1 stands for a perfect positive relationship. In Figure 12(b), all four curves start from similar Y values but end at distinctive Y values, where the LipidBERT model achieves significantly higher values (0.98 and 0.97) than the AGILE model (0.37 and 0.23) for the Hela and Raw 264.7 datasets, respectively. $R^{2}$ is perhaps the most direct measure of a predictive regression model, where higher $R^{2}$ indicates better prediction accuracy and $R^{2}$=1 denotes perfect prediction. During fine-tuning, our LipidBERT model gradually improves its $R^{2}$ and achieves 0.95 after 100 epochs. On the other hand, the AGILE model struggles to enhance its $R^{2}$ values and fluctuates around 0. To summarize, our LipidBERT model not only exhibits state-of-the-art performance on our own experimental datasets but also on the publicly available datasets.
@@ 3 Conclusion
In this study, we have introduced LipidBERT, a self-supervised pre-trained language model specifically designed for in silico generated virtual lipid molecules. Through the generation of a comprehensive database of 10 million virtual lipids using fragment-based generative methods and reinforcement learning, we have addressed the scarcity of reported lipid structures available for effective NLP pre-training. Our results demonstrate that LipidBERT, when fine-tuned for downstream lipid nanoparticle (LNP) property prediction tasks, exhibits exceptional performance with an $R^{2}$ score exceeding 0.9. This underscores the efficacy of using pre-trained language models on virtual datasets for molecular property prediction.
@@ Acknowledgements
We are grateful to the METiS platform technology team for providing wet-lab experimental data. We also extend our thanks to Shaoli Liu, Yu Lu, Ruilu Feng, and Liu Yang for insightful discussions on LNP formulation. Additionally, We appreciate the valuable input from Per Larsson (Uppsala University), Michelle Lynn Hall (Eli Lilly), and Qun Zeng (XtalPi) on the applications of AI, quantum mechanics and molecular dynamics on lipid and LNP research. We thank Xin Feng for helping construct the schematic figures. We also thank Le Yin and Fei Han for useful discussions on AI and computational lipid design, as well as Wei Xu, Hongya Han, and Beibei Cao for useful discussions on biology-related topics. Our gratitude extends to Zixuan Han, Khaled Zemoura, Xiaoyun Ma, Fang Lao, Liujun Song, Kaiqi Sun, and Yang Chen for useful discussions on multiple topics. Finally, we thank Jeff Warrington for his valuable insights on the lipid library.

### v3

@@ 1 Introduction
In this paper, we present LipidBERT, a model pre-trained on an in silico generated and meticulously screened lipid library constructed via fragment-based generative methods and reinforcement learning (RL). Molecular fragments with both lipid and small molecule features were generated by RL and connected to form ionizable lipids. These were then subjected to a series of knowledge-, molecular dynamics-, and AI-based filters. The resulting library consists of 10 million ionizable lipids, which were subsequently used for pre-training several BERT-like [21] models on different data scales. The models were then fine-tuned for downstream LNP property prediction tasks, using wet-lab labeled data from our experimental team. The fine-tuned prediction model exhibits a remarkably high Pearson correlation coefficient ( 0.8), demonstrating the power of pre-trained lipid language models on virtual datasets. Apart from language models, we also have graph-based and several distribution-based models, which might be reported in the near future.
@@ 2.3.3 Connecting Atom Prediction - Sequence Classification
Figure 5: Visualization of the predicted connecting atom between head and tail via sequence classification, using the connecting points prediction head in the pre-trained model. The predicted atoms are explicitly marked. The visualization was created using RDKit [22]. We intentionally selected six lipids with low AI-predicted values but high diversity from the sampled set.
@@ 2.4 Fine-Tuning
For each pre-trained model, we use the 768-dimensional embeddings of the [CLS] token to represent the entire sequence. The sequence output [CLS] 768-dimensional embeddings pass through a linear layer and a tanh activation function, followed by an MLP layer to generate the final output. The MLP layer consists of four layers that gradually transform the embeddings from 768 to 128 dimensions, applying the GeLU [26] activation function. A final linear layer is employed to generate the output for regression tasks. In the context below, LipidBERT refers to the model pre-trained with or without secondary tasks that achieves the best performance (highest validation Pearson correlation coefficient value). To guard against overfitting, we applied early stopping: we recorded the maximum validation Pearson correlation coefficient (PCC) obtained over 100 training epochs. The PCC ranges from -1 to 1, where -1 represents a perfect negative linear relationship and 1 stands for a perfect positive relationship.
@@ 2.4.1 Fine-Tuning Using Our Wet-Lab Experimental Dataset
The up-to-date (Mar 2025) wet-lab data used for fine-tuning were provided by our experimental team and encompass a range of lipid nanoparticle (LNP) properties. Each of our pre-trained models—originally trained on datasets of varying size—was subsequently fine-tuned on these experimental LNP measurements. Below, we present representative fine-tuning results for selected LNP properties. All data in this study were generated using fixed component ratios in the LNP formulations. The primary goal of these experiments was to screen and identify the most promising ionizable lipids for downstream development of cell-, tissue-, or organ-targeted LNPs. In this manuscript, we focus on Ex-vivo organ fluorescence intensity values; for detailed results from other property models, please contact the corresponding authors.
Figure 9: Validation set Pearson Correlation Coefficient (PCC) values change along epochs during LipidBERT fine-tuning, concerning various LNP properties.
Organ Ex-vivo Fluorescence Intensity Prediction
First, we evaluated LipidBERT on downstream sequence-to-property regression tasks using up-to-date wet-lab fluorescence intensity (FI) measurements in five organs (heart, liver, spleen, lung, and kidney as given in Figure 9). Our model achieved excellent validation Pearson correlation coefficients (PCC), all approaching or exceeding 0.80. In particular, liver and spleen converged most rapidly, stabilizing around 0.87, while the lung FI curve displayed greater variability. It is important to note that LipidBERT performs especially well on ionizable lipids whose scaffolds are well represented in our experimental set; its predictions may be less accurate for compounds with entirely novel scaffolds. As we continue to expand both the size and structural diversity of our lipid library, we expect the model’s predictive accuracy and generalizability to improve further.
Figure 10: Validation set Pearson Correlation Coefficient (PCC) values change along epochs concerning (a) ex-vivo fluorescence intensity in liver, and (b) spleen. Different curves represent LipidBERT models pre-trained on different scales of training dataset.
@@
The effect of training set sizes is also explored by fine-tuning LipidBERT models pre-trained on various training set sizes using the same experimental data mentioned above. The variation of the validation PCC values over training epochs is presented in Figure 10. Maximum validation PCC values are selected to evaluate the performance of each model. From both figures, it is evident that model performance improves as the size of the training dataset increases. In Figure 10(a), there is a dramatic enhancement in PCC from 0.37 with the 0.05M training set to 0.64 with the 0.25M training set, and further to 0.75 with the 1M training set. Increasing the training dataset size from 1M to 5M boosts the PCC to 0.80, while expanding the dataset from 5M to 10M yields a increase, with PCC reaching 0.87. In contrast, Figure 10(b) shows that the 0.05M and 0.25M models exhibit negligible differences in predicting Ex-vivo spleen FI, with PCC values of 0.31 and 0.30, respectively. However, there is a significant increase in PCC between the 0.25M and 1M models, where the PCC jumps from 0.30 to 0.67. Further increases in dataset size eventually yield a 10M model with an PCC value of 0.87. These examples demonstrate that increasing the training dataset size is critical to obtaining a fine-tuned model with high PCC values, highlighting the importance of large datasets for model accuracy.
Figure 11: Validation set Pearson Correlation Coefficient (PCC) values of LipidBERT and various XGBoost models trained on combinations of different descriptors.
@@
We also compared the fine-tuned 10M-LipidBERT model with our previously developed XGBoost [27] model, which is trained using combinations of regular molecular descriptors, Molecular Dynamics (MD) descriptors, and Quantum Mechanics (QM) descriptors. The regular descriptors are generated via RDKit [22], while the MD and QM descriptors are generated from MD and QM simulations, respectively. The MD descriptors are obtained from our fully-automated high-throughput (HTS) lipid bilayer simulations, designed to mimic the lipid movement in realistic LNPs. Given a lipid SMILES, the HTS platform can automatically construct the lipid/bilayer 3D structures, perform MD simulations, and extract the MD descriptors from the MD trajectories without human intervention. Similarly, the QM descriptors are also obtained automatically from our HTS platform. QM calculations are performed on a single ionizable lipid molecule to extract the QM descriptors. All descriptors are combined and used for XGBoost training on the overall fluorescence intensity, and the results are presented in Figure 11. The figure shows that the LipidBERT model exhibits overwhelming advantages over all XGBoost models, achieving an exceptionally high PCC value of 0.87. Among the XGBoost models, the QM + regular descriptors-based model gives the highest PCC value of 0.62, indicating that the combination of QM simulations and regular RDKit descriptors can derive better descriptors than other cases.
@@
Apart from LipidBERT, we also trained PhatGPT (Precision High-throughput AI-driven Targeting using Generative Pre-trained Transformer), our GPT-like [28] lipid generation model using the same dataset as LipidBERT. Despite the fact that PhatGPT is originally designed and trained for generative tasks, we compared the performance of the two language models on downstream sequence regression tasks. It is important to note that the vanilla PhatGPT has different model parameters (such as the number of attention layers and embedding sizes). For the purpose of comparison, we pre-trained another PhatGPT model using the same parameters as LipidBERT. Several LNP properties were used for fine-tuning, but PhatGPT did not perform as expected, yielding validation PCC values around 0.3. This outcome further demonstrates the superiority of BERT-like models in sequence regression/classification tasks, while GPT-like models are primarily designed for token generation.
Diligent
Diligent (Diffusion-like Lipid Generation Technology) is our diffusion-based generative framework trained on a dataset of 10 million lipid structures, realized in both sequence- and graph-based architectures. The model leverages a DiT [29] backbone, in which a transformer encoder is trained to predict the Gaussian noise at each diffusion step. In contrast to image-based diffusion—where inputs are first partitioned into patches for pixel-level processing—Diligent operates directly on molecular representations, diffusing either SMILES tokens or graph nodes through the forward and reverse processes. Comprehensive evaluation of these lipid generation models will be presented in a forthcoming publication.
@@ 2.4.2 Fine-Tuning Using the AGILE Public Dataset
We followed the official instructions on AGILE’s github page (https://github.com/bowang-lab/AGILE) to fine-tune the pre-trained AGILE model utilizing its provided dataset, and compared it with our fine-tuned LipidBERT model using the same dataset. Both models were fine-tuned for 100 epochs using the same random seed for an 80%-10%-10% train-validation-test dataset split. The fine-tuned results, including validation loss and Pearson correlation coefficient (PCC), are given in Figure 12. Generally speaking, our transformer-based LipidBERT model outperforms the graph-based AGILE model in most evaluation metrics, except for the Raw 264.7 validation loss (2.12 vs. 2.03 for AGILE and LipidBERT, respectively). For validation loss (Figure 12(a)), LipidBERT experiences a rapid decline from 23 to 5.01 in the first 7 epochs, and then gradually drops to 2.07 for the Hela dataset, which is lower than the AGILE validation losses (12.8). It should be noted that the deviations between the two models concerning Raw 264.7 are negligible. In Figure 12(b), all four curves start from similar Y values but end at distinctive Y values, where the LipidBERT model achieves higher values (0.98 and 0.78) than the AGILE model (0.37 and 0.23) for the Hela and Raw 264.7 datasets, respectively. Despite comparable validation losses on Raw 264.7, LipidBERT demonstrates a significantly higher correlation with ground truth. To summarize, our LipidBERT model not only exhibits state-of-the-art performance on our own experimental datasets but also on the publicly available datasets.
@@ 3 Conclusion
In this study, we have introduced LipidBERT, a self-supervised pre-trained language model specifically designed for in silico generated virtual lipid molecules. Through the generation of a comprehensive database of 10 million virtual lipids using fragment-based generative methods and reinforcement learning, we have addressed the scarcity of reported lipid structures available for effective NLP pre-training. Our results demonstrate that LipidBERT, when fine-tuned for downstream lipid nanoparticle (LNP) property prediction tasks, exhibits exceptional performance with Pearson correlation coefficient values  0.8. This underscores the efficacy of using pre-trained language models on virtual datasets for molecular property prediction.
@@ Acknowledgements
We are grateful to the METiS platform technology team for providing wet-lab experimental data. We also extend our thanks to Hongming Chen, Shaoli Liu, Yu Lu, Ruilu Feng, and Liu Yang for insightful discussions on LNP formulation. Additionally, we appreciate the valuable input from Per Larsson (Uppsala University), Michelle Lynn Hall (Eli Lilly), and Qun Zeng (XtalPi) on the applications of AI, quantum mechanics and molecular dynamics on lipid and LNP research. We thank Xin Feng for helping construct the schematic figures. We also thank Le Yin and Fei Han for useful discussions on AI and computational lipid design, as well as Wei Xu, Hongya Han, and Beibei Cao for useful discussions on biology-related topics. Our gratitude extends to Zixuan Han, Khaled Zemoura, Xiaoyun Ma, Fang Lao, Liujun Song, Kaiqi Sun, and Yang Chen for useful discussions on multiple topics. Finally, we thank Jeff Warrington for his valuable insights on the lipid library.